### v2

@@ 1 Introduction
Reinforcement learning (RL) represents a significant paradigm [120] within machine learning (ML), focused on optimizing decision processes through interaction with an environment. In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks [94].
Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [107, 85, 145]. In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent. The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios. These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. Self-play provides an approach where an agent interacts with copies or past versions of itself [108, 5]. This method promises a more stable and manageable learning process. The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go [115, 117, 116, 112], chess [116, 112], poker [95, 50], and video games [7, 129]. In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements[115, 116].
Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)[134], relatively few comprehensive surveys focus solely on self-play algorithms. Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms [54], while a separate study focuses exclusively on PSRO [8], without considering other self-play algorithms. Another study examines the theoretical safety of self-play [34]. While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play. Therefore, this survey aims to fill this gap.
@@ 2.1 RL Framework
In RL, the environment is modeled as an MDP, where the Markovian assumption states that the environment’s evolution is fully determined by its current state, eliminating the need to consider past states. An agent interacts with the environment by taking actions, which result in different states and associated rewards. MDPs can be extended to multi-agent settings, known as Markov games (MGs) [79] or stochastic games [113]. We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, meaning they have limited access to the environment’s state.
A POMG $\mathcal{G}$ is defined by $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma,\rho)$. $\mathcal{N}=\{1,\cdots,n\}$ denotes n agents. $\mathcal{S}$ is the state space. $\mathcal{A}=\prod_{i=1}^{n}\mathcal{A}_{i}$ is the product of the action space of each agent. Similarly, $\mathcal{O}=\prod_{i=1}^{n}\mathcal{O}_{i}$ is the product of the observation space of each agent. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability from one state to another given the actions of each agent. =$\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$. $\gamma\in[0,1]$ is the discount factor. $\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution. If it is a fully cooperative setting, agents can share the same reward function [145, 106, 39, 68]. When $n=1,\mathcal{O}_{i}=\mathcal{S}$, the POMG returns to the MDP.
@@ 2.2.1 Perfect Information and Imperfect Information
In a game with perfect information, only one player moves at a time. Each player comprehensively understands the current game state, the complete history of moves made, and the potential future outcomes. If these conditions are not met, the game is considered to have imperfect information [84, 99]. An example of a game with perfect information is Go. In Go, both players have full access to the game state at all times, including the positions of all stones and the identity of the next player. In contrast, Texas Hold’em poker is a game with imperfect information. Players cannot see their opponents’ private cards, creating uncertainty and requiring decisions based on probabilistic reasoning, bluffing, and interpreting others’ actions.
@@ 2.2.3 Transitive Game and Non-transitive Game
For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes follow a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists\leavevmode\nobreak\ \pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A classical example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
@@ 2.2.5 Nash Equilibrium
However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank [101] and Correlated Equilibrium (CE) [2] as alternatives. Some studies also resort to Replicator Dynamics [124] to analyze strategy evolution.
@@ 3 Algorithms
Based on existing self-play work [54, 73, 82, 42], we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. For simplicity, our framework is illustrated for symmetric games. All players share a policy population with a fixed maximum size. In each iteration, a newly initialized policy is trained, while opponent policies are sampled from the population. After training, the new policy is added to the population. The updated policy population is then evaluated, and then the opponent sampling strategy is recalculated for the next iteration.
@@ 3.1 Framework Definition
Figure 3: Illustration of our framework. The top row shows two different initialization methods for the policy population $\Pi$ and the interaction matrix $\Sigma$: lazy initialization and immediate initialization (corresponding to Line 1 in Algo. 1). The bottom row illustrates two training paradigms: standard training and ongoing training (corresponding to Lines 2~9 in Algo. 1). We categorize self-play algorithms into four types: traditional self-play, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Among these, traditional self-play, the PSRO series, and the regret-minimization-based series utilize lazy initialization and standard training, whereas the ongoing-training-based series employs immediate initialization and ongoing training. For a more detailed description and analysis, please refer to Sec.3 and Table1.
Algorithm 1 the Framework of Self-play.
@@
In Algo. 1, we define a unified self-play framework based on [82, 73, 42, 54]. We combine visual illustrations (Fig. 3) with detailed textual descriptions to explain the key processes of our framework. Next, we provide an in-depth explanation of these processes:
Figure 4: Interaction matrix $\Sigma$ examples. When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs. Here, we consider three representative self-play algorithms. In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$. Note that this figure is reproduced from [82] and this concept is initially proposed by [42].
  • $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which provides supplementary information for certain algorithms. We denote the policy $\pi_{i}(\cdot|h(i))$ as $\pi_{i}^{h}$. Note that $i$ refers to the $i$th policy in the population, not the $i$th player, and $N$ denotes the policy population size, not the number of players.
    $\Pi$ can be initialized (Line 1 in Algo. 1) in two ways: lazy initialization and immediate initialization. In lazy initialization, $\Pi$ starts with $N$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo. 1). In immediate initialization, $\Pi$ is initialized upfront with $N$ real policies, which may be randomly generated or pre-trained models. A policy $\pi_{i}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{i}$, where each $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$, indicating how to sample opponents’ policies. The dimension $C_{1}$ varies according to the sampling method. For instance, $\sigma_{i}$ can represent the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=N^{n-1}$ and $n$ denotes the number of players. Alternatively, $\sigma_{i}$ can be viewed as parameters of a sampling network. Specially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 4).
    $\Sigma$ will be initialized (Line 1 in Algo. 1) based on the initialization of $\Pi$. If $\Pi$ uses lazy initialization, $\Sigma$ will be initialized as an empty matrix. If $\Pi$ uses immediate initialization, $\Sigma$ will be calculated based on $\Pi$. We will discuss this in more detail later.
@@
  • Initialize $\pi_{i}^{h}$ (Line 4 in Algo. 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used. It can be initialized randomly or by leveraging pre-trained models [115, 129] or through a recently updated policy.
  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 5 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE [92]. However, it often requires considerable computational effort. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) [96, 86] or introduced to enhance diversity [4, 103, 83].
    Some details of Algo. 23 and 4 need to be mentioned. Algo. 2 is a RL method that can be widely used across different categories. Moreover, off-policy RL algorithms typically require a replay buffer to gather samples; however, they are not explicitly included for simplicity. Algo. 3 is specifically utilized by Regularized Nash Dynamics (R-NaD) [105] and Algo. 4 is specifically utilized by the regret-minimization-based series. They will be introduced in detail in Sec. 3.3.5 and Sec. 3.5 respectively.
@@
  • $\textit{MSS}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$ (Line 7 in Algo. 1): A meta-strategy solver (MSS) MSS takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output. If the MSS produces a constant matrix irrespective of the input, the evaluation step (Line 6 in Algo. 1) can be skipped to reduce computations.
    Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo. 1) by first evaluating the performance and then applying the MSS.
@@
4:     Use regret matching to play the game and obtain new regret minimization information added to $h(i)$.

### v3

@@ 1 Introduction
Reinforcement learning (RL) represents a significant paradigm (Sutton & Barto, 2018) within machine learning (ML), focused on optimizing decision processes through interaction with an environment. In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks (Mnih et al., 2013).
Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics (Rashid et al., 2018; Mahajan et al., 2019; Yu et al., 2022). In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent. The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios. These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. Self-play provides an approach where an agent interacts with copies or past versions of itself (Samuel, 2000; Bansal et al., 2018). This method promises a more stable and manageable learning process. The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go (Silver et al., 2016; 2017; 2018; Schrittwieser et al., 2020), chess (Silver et al., 2018; Schrittwieser et al., 2020), poker (Moravčík et al., 2017; Heinrich et al., 2015), and video games (Berner et al., 2019; Vinyals et al., 2019). In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements(Silver et al., 2016; 2018).
Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)(Wellman et al., 2025), relatively few comprehensive surveys focus solely on self-play algorithms. Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms (Hernandez et al., 2021), while a separate study focuses exclusively on PSRO (Bighashdel et al., 2024), without considering other self-play algorithms. Another study examines the theoretical safety of self-play (DiGiovanni & Zell, 2021). While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play. Therefore, this survey aims to fill this gap.
@@ 2.1 RL Framework
In RL, the environment is modeled as an MDP, where the Markovian assumption states that the environment’s evolution is fully determined by its current state, eliminating the need to consider past states. An agent interacts with the environment by taking actions, which result in different states and associated rewards. MDPs can be extended to multi-agent settings, known as Markov games (MGs) (Littman, 1994) or stochastic games (Shapley, 1953). We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, meaning they have limited access to the environment’s state.
A POMG $\mathcal{G}$ is defined by $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma,\rho)$. $\mathcal{N}=\{1,\cdots,n\}$ denotes n agents. $\mathcal{S}$ is the state space. $\mathcal{A}=\prod_{i=1}^{n}\mathcal{A}_{i}$ is the product of the action space of each agent. Similarly, $\mathcal{O}=\prod_{i=1}^{n}\mathcal{O}_{i}$ is the product of the observation space of each agent. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability from one state to another given the actions of each agent. =$\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$. $\gamma\in[0,1]$ is the discount factor. $\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution. If it is a fully cooperative setting, agents can share the same reward function (Yu et al., 2022; Rashid et al., 2020; Foerster et al., 2018; Kuba et al., 2021). When $n=1,\mathcal{O}_{i}=\mathcal{S}$, the POMG returns to the MDP.
@@ 2.2.1 Perfect Information and Imperfect Information
In a game with perfect information, only one player moves at a time. Each player comprehensively understands the current game state, the complete history of moves made, and the potential future outcomes. If these conditions are not met, the game is considered to have imperfect information (Lucchetti, 2011; Mycielski, 1992). An example of a game with perfect information is Go. In Go, both players have full access to the game state at all times, including the positions of all stones and the identity of the next player. In contrast, Texas Hold’em poker is a game with imperfect information. Players cannot see their opponents’ private cards, creating uncertainty and requiring decisions based on probabilistic reasoning, bluffing, and interpreting others’ actions.
@@ 2.2.3 Transitive Game and Non-transitive Game
For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes follow a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists~{}\pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A classical example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
@@ 2.2.5 Nash Equilibrium
However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank (Omidshafiei et al., 2019) and Correlated Equilibrium (CE) (Aumann, 1974) as alternatives. Some studies also resort to Replicator Dynamics (Taylor & Jonker, 1978) to analyze strategy evolution.
@@ 3 Algorithms
Based on existing self-play work (Hernandez et al., 2021; Lanctot et al., 2017; Liu et al., 2022c; Garnelo et al., 2021), we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. For simplicity, our framework is illustrated for symmetric games. All players share a policy population with a fixed maximum size. In each iteration, a newly initialized policy is trained, while opponent policies are sampled from the population. After training, the new policy is added to the population. The updated policy population is then evaluated, and then the opponent sampling strategy is recalculated for the next iteration.
@@ 3.1 Framework Definition
Figure 3: Illustration of our framework. The top row shows two different initialization methods for the policy population $\Pi$ and the interaction matrix $\Sigma$: lazy initialization and immediate initialization (corresponding to Line 1 in Algo. 1). The bottom row illustrates two training paradigms: standard training and ongoing training (corresponding to Lines 2~9 in Algo. 1). We categorize self-play algorithms into four types: traditional self-play, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Among these, traditional self-play, the PSRO series, and the regret-minimization-based series utilize lazy initialization and standard training, whereas the ongoing-training-based series employs immediate initialization and ongoing training. For a detailed description and analysis, please refer to Sec.3 and Table1.
Algorithm 1 A unified framework of self-play.
@@
In Algo. 1, we define a unified self-play framework based on Liu et al. (2022c); Lanctot et al. (2017); Garnelo et al. (2021); Hernandez et al. (2021). We combine visual illustrations (Fig. 3) with detailed textual descriptions to explain the key processes of our framework. Next, we provide an in-depth explanation of these processes:
  • $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which provides supplementary information for certain algorithms. We denote the policy $\pi_{i}(\cdot|h(i))$ as $\pi_{i}^{h}$. Note that $i$ refers to the $i$th policy in the population, not the $i$th player, and $N$ denotes the policy population size, not the number of players. $\Pi$ can be initialized (Line 1 in Algo. 1) in two ways: lazy initialization and immediate initialization. In lazy initialization, $\Pi$ starts with $N$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo. 1). In immediate initialization, $\Pi$ is initialized upfront with $N$ real policies, which may be randomly generated or pre-trained models. A policy $\pi_{i}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{i}$, where each $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$, indicating how to sample opponents’ policies. The dimension $C_{1}$ varies according to the sampling method. For instance, $\sigma_{i}$ can represent the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=N^{n-1}$ and $n$ denotes the number of players. Alternatively, $\sigma_{i}$ can be viewed as parameters of a sampling network. Specially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 4). $\Sigma$ will be initialized (Line 1 in Algo. 1) based on the initialization of $\Pi$. If $\Pi$ uses lazy initialization, $\Sigma$ will be initialized as an empty matrix. If $\Pi$ uses immediate initialization, $\Sigma$ will be calculated based on $\Pi$. We will discuss this in more detail later.
@@
  • Initialize $\pi_{i}^{h}$ (Line 4 in Algo. 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used. It can be initialized randomly or by leveraging pre-trained models (Silver et al., 2016; Vinyals et al., 2019) or through a recently updated policy.
  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 5 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE (McMahan et al., 2003). However, it often requires considerable computational effort. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) (Muller et al., 2020; Marris et al., 2021) or introduced to enhance diversity (Balduzzi et al., 2019; Perez-Nieves et al., 2021; Liu et al., 2021). Some details of Algo. 23 and 4 need to be mentioned. Algo. 2 is a RL method that can be widely used across different categories. Moreover, off-policy RL algorithms typically require a replay buffer to gather samples; however, they are not explicitly included for simplicity. Algo. 3 is specifically utilized by Regularized Nash Dynamics (R-NaD) (Perolat et al., 2021) and Algo. 4 is specifically utilized by the regret-minimization-based series. They will be introduced in detail in Sec. 3.3.5 and Sec. 3.5 respectively.
@@
  • $\textit{MSS}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$ (Line 7 in Algo. 1): A meta-strategy solver (MSS) MSS takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output. If the MSS produces a constant matrix irrespective of the input, the evaluation step (Line 6 in Algo. 1) can be skipped to reduce computations. Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo. 1) by first evaluating the performance and then applying the MSS.
Figure 4: Interaction matrix $\Sigma$ examples. When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs. Here, we consider three representative self-play algorithms. In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$. Note that this figure is reproduced from Liu et al. (2022c) and this concept is initially proposed by Garnelo et al. (2021).
@@
4:     Use regret matching to play the game and update $h(i)$ with new regret minimization information.